# MUG: Pseudo Labeling Augmented Audio-Visual Mamba Network for Audio-Visual Video Parsing

### v1

@@ Abstract
The weakly-supervised audio-visual video parsing (AVVP) aims to predict all modality-specific events and locate their temporal boundaries. Despite significant progress, due to the limitations of the weakly-supervised and the deficiencies of the model architecture, existing methods are lacking in simultaneously improving both the segment-level prediction and the event-level prediction. In this work, we propose a audio-visual Mamba network with pseudo labeling aUGmentation (MUG) for emphasising the uniqueness of each segment and excluding the noise interference from the alternate modalities. Specifically, we annotate some of the pseudo-labels based on previous work. Using unimodal pseudo-labels, we perform cross-modal random combinations to generate new data, which can enhance the model’s ability to parse various segment-level event combinations. For feature processing and interaction, we employ a audio-visual mamba network. The AV-Mamba enhances the ability to perceive different segments and excludes additional modal noise while sharing similar modal information. Our extensive experiments demonstrate that MUG improves state-of-the-art results on LLP dataset in all metrics ( e.g., gains of 2.1% and 1.2% in terms of visual Segment-level and audio Segment-level metrics). Our code is available at https://github.com/WangLY136/MUG.
@@ 1 Introduction
Fig.1: (a): With only video level label, AVVP requires parsing out visual events, audio events, and their temporal boundaries. (b): Constrained by the weakly-supervision learning and the noise in LLP dataset, the model fails to learn a large number of event combinations at the segment-level, and a large quantity of empty pseudo-labels occur. Meanwhile, the input tokens are approaching the long sequence threshold of ViT-S [45].
Limited by weakly-supervised learning, the model can only learn the features of different events from video-level labels. The model fails to learn a large number of event combinations at the segment-level, which significantly affects the model’s prediction of segment-level events. Previous [23, 9] works have been devoted to extracting pseudo-labels. However, due to the presence of noise in the labels of the LLP dataset, the quality of the pseudo-labels is restricted. Meanwhile, the events occurring in each modality are independent of each other and may even be completely unrelated. An event in one modality may either complement the prediction of another modality or introduce noise. Some works [8, 44, 29, 35] attempt to develop more robust audio-visual encoders for embedding more effective audiovisual features, but they are insufficient in retaining the original modality features while sharing the features. Finally, the AVVP task requires the simultaneous prediction of events at both the Segment-level (single frame) and the Event-level (multiple frames). Therefore, it is necessary not only to conduct inference on single-frame images/audio but also to capture the strong causal relationships among multiple frames of images/audio. The commonly used Transformer (primarily due to its attention mechanism) models in the past have certain limitations when dealing with long sequences (a large number of tokens). The token length of the visual input sequence in AVVP is close to 2000, which is approaching the threshold of ViT-S [45]. Recently, some Mamba-based backbones [13, 6] have demonstrated great potential in long-sequence modeling and can utilize the causal order to model sequences. However, for single-frame image, there is no sequential dependency. Instead, the ability to model the overall space is required [45, 17]. Therefore, Mamba has deficiencies in the recognition of single-frame image.
@@
Inspired by the Mamba-Transformer architecture in previous works [17, 25, 11], based on the HAN [34] model, we propose a brand-new baseline that can simultaneously improve the model’s performance at both the segment-level and the event-level. Specifically, we first utilize the spatial-temporal attention composed of Mamba to capture the key information in the sequence. Subsequently, we propose a cross-modal adaptive mamba fusion structure. This structure captures the cross-modal similar information while retaining the intermodal information through a shared matrix. In order to prevent the causal model from forgetting the early token information, we add an additional dynamic branch to alleviate this problem. To enhance the cross-modal similar information captured in the previous step, we introduce a Mamba feature enhancement module. We incorporate the HAN model at the end of the network, and its simple Transformer (attention architecture) further strengthens the long-range spatial dependencies.
@@ 2 Related Works
State Space Models (SSM). The state space model [15] has become one of the most important backbones in deep learning. It originates from classical control theory and provides linear scalability of sequence length for modeling remote dependency. To enhance the practical feasibility, Gu et al. [14] proposed the S4 module. Subsequently, Smith et al. [32] proposed a SSM model supporting multiple inputs and multiple outputs, and Hasani et al. [16] proposed a SSM model for liquid structures. Based on this, Gu et al. [13] proposed the mamba architecture. It merges the previous SSM structure with the MLP module in Transformer into a single module, thus obtaining an architectural design with a selective state space. Inspired by ViT [7] and Swin Transformer [27], Zhu et al. [50] and Liu et al. [26] proposed Vision Mamba and VMamba, respectively, which showed impressive potential on numerous vision tasks. Meanwhile, Mamba structures have begun to emerge on multimodal tasks. xie et al. [21] proposed a mamba multimodal fusion structure for medical images. Li et al. [24] utilised a coupled state space model to enhance information fusion between different modalities, which greatly improves the inference speed. All these works laid the foundation for subsequent multimodal mamba research.
@@ 3 Our Approach
We make improvements from both the data and model perspectives. On the data side, we propose a data augmentation method specifically for the AVVP task. We first annotate some obviously erroneous pseudo-labels. Building on this, we randomly combine the visual and audio modalities of any two videos to generate new data. This method effectively enhance the quality of the dataset, allowing the model to learn the features of each segment more thoroughly. On the model side, we propose a audio-visual mamba network. We capture both temporal and local information simultaneously using temporal-spatial attention and share some parameters of the SSM. This approach not only enhances the model’s perception of each segment but also shares cross-modal similar information while preserving unimodal information. For effective projection and better model optimization, we incorporate the segment-wise pseudo labels generated in recent work [46] to provide fine-grained supervision. We also introduce the text modality to exclude irrelevant modality noise. Our approach MUG framework is shown in Fig.3. We describe the problem formulation in section 3.1. Then we illustrate the data augmentation and mamba framework in section 3.2 and section 3.3.
@@ 3.2 Data augmentation in AVVP
Manual annotation (MA). Lai et al.[23] used frozen CLIP and CLAP to compute the visual segment pseudo-labels $\hat{y}^{m}_{v}$ and audio segment pseudo-labels $\hat{y}^{m}_{a}$, respectively. In process of generating pseudo-labels, Lai et al. use the pre-trained CLIP and CLAP to compute the cosine similarity of the features and the 25 categories of labels, and pass a threshold to obtain one-hot encoded pseudo-labels. However, the LLP dataset itself is not annotated sufficient accuracy and the 25 categories cannot encompass all events. The limitation of the LLP dataset is one of the reasons that lead to the insufficient accuracy of pseudo-labels. We convert the generated one-hot pseudo-labels to CSV file, and we can find that there are many null labels (i.e., no events occurred) in the visual modality, which is abnormal since a video may lack sound but rarely lacks imagery. For such pseudo-labels, we compare them with the LLP dataset and manually annotate them. Pseudo-labels that belong to the 25 categories are accurately annotated, while those that do not belong to one of the 25 categories will not be annotated and will be discarded in cross modality random combination. More annotation details and ablation experiments are shown in the supplementary.
Cross modality random combination (CMRC). The inability to use data enhancement in previous work was due to the absence of unimodal label information. With the availability of high-quality pseudo-labels, data enhancement for the AVVP task became feasible. For the LLP dataset, we first count the distribution of all video-level labels that appear more than 50 times in the entire training set. Based on this distribution, we selectively combine the visual features of one video with the audio features of another video to generate new video data. The label of the new video is the union of the visual modality pseudo-labels and the audio modality pseudo-labels. Fig.2 represents the process of the CMRC. The pseudo-labels that cannot be annotated and their corresponding videos (i.e., pseudo-labels and videos that do not fall within the 25 categories) are eliminated when generating new data. In order to avoid overfitting and introduce excessive additional data noise, we generate five batches of data according to the actual distribution, with quantities of 1585, 3242, 4610, 6080 and 12096 respectively. We separately test data augmentation for different batches to find the critical point of this method. When the amount of generated data is small, there is still room for model optimization. When the generated data is abundant, the noise in the dataset will increase, and meanwhile, the risk of overfitting rises. CMRC is justified because in the real world, visual and audio modalities are not always correlated, and sound signals can originate from various directions within a location. More details of the data distribution and combination process are presented in the supplementary.
@@ 3.3 AV-Mamba Network
Temporal-Spatial attention (TSA). We extract audio and visual features by using pre-trained VGGish and ResNet-152, denoted as $\{f^{a}_{t}\}^{T}_{t=1}$, $\{f^{v}_{t}\}^{T}_{t=1}$. $F^{a}$ and $F^{v}$ stand for the feature set in the same video, which are defined as $F^{a}=\{f^{a}_{1},...,f^{a}_{T}\}\in\mathbb{R}^{T\times d}$ and $F^{v}=\{f^{v}_{1},...,f^{v}_{T}\}\in\mathbb{R}^{T\times d}$, d is the feature dimension. To aggregate features from different segments and enhance the feature expressiveness of each segment, we propose a segment-based temporal-spatial attention module, analogous to the convolutional block attention module [36], but implement it using Mamba.
For each segment, we generate two feature vectors by global maximum pooling and global average pooling. After the feature vectors are fed into a shared fully connected layer, we can obtain the final attention weight vector $W^{m}_{t}$:
@@
where $f^{m}_{t}$ denotes audio or visual features within a video, whose dimensions are $(B,T,D)$, representing batchsize, segments, and feture, respectively. $\delta$ indicates sigmoid function, $m\in\{a,v\}$
Similar to temporal attention, we generate two feature vectors by max pooling and average pooling along the temporal (segment) dimension. The features derived from max pooling and average pooling are then concatenated along the temporal dimension to obtain a feature vector, which is referred to as spatial attention $S^{m}_{t}$, $m\in\{a,v\}$ :
@@
The outputs of the channel attention module and the spatial attention module are element-wise multiplied to obtain the final attention-enhanced feature $\hat{f}^{m}_{t}$, $m\in\{a,v\}$:
@@
Adaptive mamba fusion (AMF). In previous work [42, 40, 28], there have been explorations in multimodal Mamba fusion with surprising results. Building upon these findings, we propose an adaptive mamba fusion network for selective interaction of visual and audio features. Specifically, after temporal-spatial attention, visual features and audio features are fed into AMF. Each of the two input features in the AMF is processed through four distinct branches, which capture unique features. $f^{Forward}_{m}$ is forward process, performing forward scans. $f^{Backward}_{m}$ is backward process, performing backward scans. The forward and backward scanning can be expressed as:
$f^{Forward}_{m}=SSM([\hat{f}^{m}_{1},\hat{f}^{m}_{2},...,\hat{f}^{m}_{N}],Conv1d)$ (4)
$f^{Backward}_{m}=SSM([\hat{f}^{m}_{N},\hat{f}^{m}_{N-1},...,\hat{f}^{m}_{1}],Conv1d)$ (5)
@@
where $Conv1d$ represents 1D convolution, $SSM$ represents the state space model, $m\in\{a,v\}$. However, only performing forward and backward scans can potentially ignore the first and last layer tokens. The Mamba has some risk of forgetting the initial token entered, so relying only on forward and backward is not adequate way to model all layers of features. For the input features, the importance of each segment is different and the events that occur are different. We have incorporated the dynamic scanning scheme [28] from recent work, allowing the scanning process to start and end at any layer. This approach enhances the model’s comprehension of different segments and thus improves the ability to parse segment-level events.
@@
Furthermore, a portion of the parameters are shared between the forward/backward SSM in both modalities. As shown in Fig.3, forward SSM and backward SSM include 6 parameters: input transition matrices A state transition matrices B,and output transition matrices C. The state transition matrices B has the most significant impact on the system, as they govern the evolution of the current hidden state [42]. Consequently, We share parameters of state transition matrices B between two modalities, while keeping input transition matrices a and output transition matrices C independent. This strategy captures cross-modal similarity information while maintaining unimodal independence. We employ multiple learnable parameters to control the degree of interaction between different modalities, preventing excessive interference between modalities. Subsequently, each feature of both modalities is subjected to a gating strategy in order to fuse the scanning results and obtain the feature $f^{AMF}_{m}$ by element-by-element summation, $m\in\{a,v\}$. Finally, as shown in Fig.3, the features undergo a simple interaction and concat to obtain mixed feature $f^{AMF}_{mix}$. $f^{AMF}_{m}$ and $f^{AMF}_{mix}$ are fed as three input to the Mamba feature enhancement module. Additionally, by modeling sequences in causal order, Mamba strengthens the connections between adjacent frames and enhance the perception ability of multi-frame events.
@@
Pseudo-label semantic interaction module (PLSIM). The information in text can be used as a cue to effectively improve the performance of the model[43, 49, 41, 35]. The above work has proven that text can serve as a priori information to culling out the other modal noise and constrain uni-modal event prediction. Contrary to the pseudo-labels employed in [23] , we eliminate logical operations with video-level labels when generating pseudo-labels for testing purposes. New pseudo-labels can be expressed as $\hat{y}^{m}_{t}$.
@@
where $\odot$ denotes Hadamard product. $\gamma_{a1}$ and $\rho_{v1}$ denote scale scaling, $\gamma_{a2}$ and $\rho_{v2}$ denote bias control. $F_{a}$ and $F_{v}$ are audio and visual features fused with semantic features. It is worth noting that when generating new data using CMRC, PLSIM will also to carry out expansion to achieve the fusion of the corresponding features.

### v2

@@ Abstract
The weakly-supervised audio-visual video parsing (AVVP) aims to predict all modality-specific events and locate their temporal boundaries. Despite significant progress, due to the limitations of the weakly-supervised and the deficiencies of the model architecture, existing methods are lacking in simultaneously improving both the segment-level prediction and the event-level prediction. In this work, we propose an audio-visual Mamba network with pseudo labeling aUGmentation (MUG) for emphasising the uniqueness of each segment and excluding the noise interference from the alternate modalities. Specifically, we annotate some of the pseudo-labels based on previous work. Using unimodal pseudo-labels, we perform cross-modal random combinations to generate new data, which can enhance the model’s ability to parse various segment-level event combinations. For feature processing and interaction, we employ an audio-visual mamba network. The AV-Mamba enhances the ability to perceive different segments and excludes additional modal noise while sharing similar modal information. Our extensive experiments demonstrate that MUG improves state-of-the-art results on LLP dataset in all metrics ( e.g., gains of 2.1% and 1.2% in terms of visual Segment-level and audio Segment-level metrics). Our code is available at https://github.com/WangLY136/MUG.
@@ 1 Introduction
Fig.1: (a): With only video level label, AVVP requires parsing out visual events, audio events, and their temporal boundaries. (b): Constrained by the weakly-supervision learning and the noise in LLP dataset, previous works fail to learn a large number of event combinations at the segment-level, and a large quantity of empty pseudo-labels occur. Meanwhile, the input tokens are approaching the long sequence threshold of ViT-S [45].
@@
Limited by weakly-supervised learning, the models can only learn the features of different events from video-level labels. Although previous works [23, 9] have been devoted to extracting pseudo-labels, the quality of the pseudo-labels is restricted due to the presence of noise in the labels of the LLP dataset. Existing frameworks fail to learn a large number of event combinations at the segment-level, which significantly affects the model’s prediction of segment-level events. Meanwhile, the events occurring in each modality are independent of each other and may even be completely unrelated. An event in one modality may either complement the prediction of another modality or introduce noise. Some works [8, 44, 29, 35] attempt to develop more robust audio-visual encoders for embedding more effective audiovisual features, but they are insufficient in retaining the original modality features while sharing the features. Finally, the AVVP task requires the simultaneous prediction of events at both the Segment-level (single frame) and the Event-level (multiple frames). Therefore, it is necessary not only to conduct inference on single-frame images/audio but also to capture the strong causal relationships among multiple frames of images/audio. The commonly used Transformer (primarily due to its attention mechanism) models in the past have certain limitations when dealing with long sequences (a large number of tokens). The token length of the visual input sequence in AVVP is close to 2000, which is approaching the threshold of ViT-S [45]. Recently, some Mamba-based backbones [13, 6] have demonstrated great potential in long-sequence modeling and can utilize the causal order to model sequences. However, for single-frame image, there is no sequential dependency. Instead, the ability to model the overall space is required [45, 17]. Therefore, Mamba has deficiencies in the recognition of single-frame image.
@@
Inspired by the Mamba-Transformer architecture in previous works [17, 25, 11], based on the HAN [34] model, we propose a brand-new baseline that can simultaneously improve the model’s performance at both the segment-level and the event-level. Specifically, we first utilize a Mamba-based attention to capture the key information in the sequence. Subsequently, we propose a cross-modal adaptive mamba fusion structure which can captures the cross-modal similar information while retaining the intermodal information through a shared matrix. In order to prevent the causal model from forgetting the early token information, we add an additional dynamic branch to alleviate this problem. To enhance the cross-modal similar information captured in the previous steps, we introduce a Mamba feature enhancement module. We incorporate the HAN model at the end of the network, and its simple Transformer (attention architecture) further strengthens the long-range spatial dependencies.
@@ 2 Related Works
State Space Models (SSM). The state space model [15] has become one of the most important backbones in deep learning. It originates from classical control theory and provides linear scalability of sequence length for modeling remote dependency. To enhance the practical feasibility, Gu et al. [14] proposed the S4 module. Subsequently, Smith et al. [32] proposed a SSM model supporting multiple inputs and multiple outputs, and Hasani et al. [16] proposed a SSM model for liquid structures. Based on this, Gu et al. [13] proposed the mamba architecture. It merges the previous SSM structure with the MLP module in Transformer into a single module, thus obtaining an architectural design with a selective state space. Inspired by ViT [7] and Swin Transformer [27], Zhu et al. [50] and Liu et al. [26] proposed Vision Mamba and VMamba, respectively, which showed impressive potential on numerous vision tasks. Meanwhile, Mamba structures have begun to emerge on multimodal tasks. xie et al. [21] proposed a mamba multimodal fusion structure for medical images. Li et al. [24] utilised a coupled state space model to enhance information fusion between different modalities, which greatly improves the inference speed. All these works lay the foundation for subsequent multimodal mamba research.
@@ 3 Our Approach
We make improvements from both the data and model perspectives. On the data side, we propose a data augmentation method specifically for the AVVP task. We first annotate some obviously erroneous pseudo-labels. Building on this, we randomly combine the visual and audio modalities of any two videos to generate new data. This method effectively enhance the quality of the dataset, allowing the model to learn the features of each segment more thoroughly. On the model side, we propose an audio-visual mamba network. We capture both temporal and local information simultaneously using Mamba-based attention and share some parameters of the SSM. This approach not only enhances the model’s perception of each segment but also shares cross-modal similar information while preserving unimodal information. For effective projection and better model optimization, we incorporate the segment-wise pseudo labels generated in recent work [46] to provide fine-grained supervision. We also introduce the text modality to exclude irrelevant modality noise. The MUG framework is shown in Fig.3. We describe the problem formulation in section 3.1. Then we illustrate the data augmentation and mamba framework in section 3.2 and section 3.3.
@@ 3.2 Data augmentation in AVVP
Manual annotation (MA). Lai et al.[23] used frozen CLIP and CLAP to compute the visual segment pseudo-labels $\hat{y}^{m}_{v}$ and audio segment pseudo-labels $\hat{y}^{m}_{a}$, respectively. In process of generating pseudo-labels, Lai et al. use the pre-trained CLIP and CLAP to compute the cosine similarity of the features and the 25 categories of labels, and pass a threshold to obtain one-hot encoded pseudo-labels. However, the LLP dataset itself is not annotated sufficient accuracy and the 25 categories cannot encompass all events. The limitation of the LLP dataset is one of the reasons that lead to the insufficient accuracy of pseudo-labels. We convert the generated one-hot pseudo-labels to CSV file, and we can find that there are many null labels (i.e., no events occurred) in the visual modality, which is abnormal since a video may lack sound but rarely lacks image. For such pseudo-labels, we compare them with the LLP dataset and manually annotate them. Pseudo-labels that belong to the 25 categories are accurately annotated, while those outside this range are left unannotated and excluded during the cross modality random combination process.
Cross modality random combination (CMRC). The inability to use data enhancement in previous works is due to the absence of unimodal label information. With the availability of high-quality pseudo-labels, data enhancement for the AVVP task became feasible. For the LLP dataset, we first count the distribution of all video-level labels that appear more than 50 times in the entire training set. Based on this distribution, we selectively combine the visual features of one video with the audio features of another video to generate new video data. The label of the new video is the union of the visual modality pseudo-labels and the audio modality pseudo-labels. Fig.2 represents the process of the CMRC. The pseudo-labels that cannot be annotated and their corresponding videos (i.e., pseudo-labels and videos that do not fall within the 25 categories) are eliminated when generating new data. In order to avoid overfitting and introduce excessive additional data noise, we generate five batches of data according to the actual distribution, with quantities of 1585, 3242, 4610, 6080 and 12096 respectively. We separately test data augmentation for different batches to find the critical point of this method. When the amount of generated data is small, there is still room for model optimization. When the generated data is abundant, the noise in the dataset will increase, and meanwhile, the risk of overfitting rises. CMRC is reasonable because in the real world, visual and audio modalities are not always correlated, because sound signals can originate from various directions.
@@ 3.3 AV-Mamba Network
Mamba-based attention (MBA). Audio and visual features are extracted using pre-trained VGGish and ResNet-152, denoted as $\{f^{a}_{t}\}^{T}_{t=1}$, $\{f^{v}_{t}\}^{T}_{t=1}$. $F^{a}$ and $F^{v}$ stand for the feature set in the same video, which are defined as $F^{a}=\{f^{a}_{1},...,f^{a}_{T}\}\in\mathbb{R}^{T\times d}$ and $F^{v}=\{f^{v}_{1},...,f^{v}_{T}\}\in\mathbb{R}^{T\times d}$, $d$ is the feature dimension. To aggregate features from different segments and enhance the feature expressiveness of each segment, we propose a segment-based attention module implemented via Mamba. This module is similar to the convolutional block attention module [36] but it is implemented by Mamba, thereby better handling the causal relationship.
For each segment, two feature vectors are generated by global maximum pooling and global average pooling. These feature vectors are fed into a shared fully connected layer, resulting in the final attention weight vector $W^{m}_{t}$:
@@
where $f^{m}_{t}$ denotes audio or visual features within a video, whose dimensions are $(B,T,D)$, representing batchsize, segments, and 128/2048 (audio/visual), respectively. $\delta$ indicates sigmoid function, $m\in\{a,v\}$. $Mamba$ denotes Mamba block.
Similar to aforementioned method, two feature vectors are produced through max pooling and average pooling along the temporal (segment) dimension. The features derived from max pooling and average pooling are then concatenated along the temporal dimension to obtain a feature vector $S^{m}_{t}$, $m\in\{a,v\}$:
@@
The outputs $W^{m}_{t}$ and $S^{m}_{t}$ are element-wise multiplied to obtain the final attention-enhanced feature $\hat{f}^{m}_{t}$, $m\in\{a,v\}$:
@@
Adaptive mamba fusion (AMF). In previous work [42, 40, 28], there have been explorations in multimodal Mamba fusion with surprising results. Building upon these findings, we propose an adaptive mamba fusion network for selective interaction of visual and audio features. Specifically, after Mamba-based attention, visual features and audio features are fed into AMF. Each of the two input features in the AMF is processed through four distinct branches, which capture unique features. $f^{Forward}_{m}$ is forward process, performing forward scans. $f^{Backward}_{m}$ is backward process, performing backward scans. The forward and backward scanning can be expressed as:
$f^{Forward}_{m}=SSM([\hat{f}^{m}_{1},\hat{f}^{m}_{2},...,\hat{f}^{m}_{N}],Conv1d),$ (4)
$f^{Backward}_{m}=SSM([\hat{f}^{m}_{N},\hat{f}^{m}_{N-1},...,\hat{f}^{m}_{1}],Conv1d),$ (5)
@@
where $Conv1d$ represents 1D convolution, $SSM$ represents the state space model, $m\in\{a,v\}$. However, only performing forward and backward scans can potentially ignore the first and last layer tokens. The Mamba has some risk of forgetting the initial token entered, so relying only on forward and backward is not adequate way to model all layers of features. For the input features, the importance of each segment is different and the events that occur are different. We have incorporated the dynamic scanning scheme from recent work [28], allowing the scanning process to start and end at any layer. This approach enhances the model’s comprehension of different segments and thus improves the ability to parse segment-level events.
@@
Furthermore, in the forward/backward SSM, two modalities share a portion of parameters due to the consistency in their scanning directions. As shown in Fig.3, forward SSM and backward SSM include 6 parameters: input transition matrices A state transition matrices B,and output transition matrices C. The state transition matrices B has the most significant impact on the system, as they govern the evolution of the current hidden state . Consequently, We share parameters of state transition matrices B between two modalities, while keeping input transition matrices a and output transition matrices C independent [42]. This strategy not only reduces the number of parameters and the potential risk of overfitting, but also preserves unimodal independence while capturing cross-modal similarity information. Subsequently, each feature of both modalities is subjected to a gating strategy in order to fuse the scanning results and obtain the feature $f^{AMF}_{m}$ by element-by-element summation, $m\in\{a,v\}$. Finally, as shown in Fig.3, the features undergo a simple add to obtain mixed feature $f^{AMF}_{mix}$. $f^{AMF}_{m}$ and $f^{AMF}_{mix}$ are fed as three input to the Mamba feature enhancement module. We employ multiple learnable parameters to control the degree of interaction between different modalities, preventing excessive interference between modalities. By modeling sequences in causal order, Mamba strengthens the connections between adjacent frames and enhance the perception ability of multi-frame events.
@@
Pseudo-label semantic interaction module (PLSIM). The information in text can be used as a cue to effectively improve the performance of the models [43, 49, 41, 35]. The above work has proven that text can serve as a priori information to culling out the other modal noise and constrain uni-modal event prediction. Contrary to the pseudo-labels employed in [23] , we eliminate logical operations with video-level labels when generating pseudo-labels for testing purposes. New pseudo-labels can be expressed as $\hat{y}^{m}_{t}$.
@@
where $\odot$ denotes Hadamard product. $\gamma_{a1}$ and $\rho_{v1}$ denote scale scaling, $\gamma_{a2}$ and $\rho_{v2}$ denote bias control. $F_{a}$ and $F_{v}$ are audio and visual features fused with semantic features. It is worth noting that when generating new data using CMRC, PLSIM will also to carry out expansion to achieve the fusion of the corresponding features. In the experiments, we find that when using batches augmented with 12,000 samples, the improvement of the PLSIM module decreased significantly and even produced negative effects. This may be due to the fact that excessive data augmentation introduces excessive label noise, thereby affecting the performance of the PLSIM module.